For the circuit shown in fig. Both voltmeter and ammeter are idealized, the battery has no appreciable internal resistance, and the ammeter reads 5 A.

a. What does the voltmeter V read?.

b. What is the total current I of the circuit?.

c. What is the total voltage of the battery?

d. What is the total resistance R of the circuit?

25.022

A

15.02

MW

45.012

15.022

10.022

35.022 V = ?


1
Expert's answer
2021-02-02T15:34:41-0500


d) First of all we must find total resistance:

"R_1 = 25\varOmega \space R_2 = 15\varOmega \space R_3 = 15\varOmega \space R_4 = 10\varOmega \space R_5 = 45\varOmega \space R_6 = 35\varOmega \space"

The "R_3 \space and \space R_4" resistors are connected in series "R_{3,4} = R_3+R_4 = 15+10=25\varOmega"

The "R_1, R_2 \space and \space R_{3,4}" resistors are connected parallel.

"\large\frac{1}{R_{1,2,3,4}}= \frac{1}{R_1} +\frac{1}{R_2}+\frac{1}{R_{3,4}} = \frac{1}{25}+\frac{1}{15}+\frac{1}{25}= \frac{11}{75} \to R_{1,2,3,4} = \frac{75}{11}\varOmega"

The "R_5 \space and \space R_{1,2,3,4}" are connected in series "R_{1,2,3,4,5} = R_{1,2,3,4}+R_5 = 45+\frac{75}{11} = 51\frac{9}{11}\varOmega"

"\large\frac{1}{R_t}= \frac{1}{R_{1,2,3,4,5}}+\frac{1}{R_6} = \frac{11}{560}+\frac{1}{35}=\frac{27}{560} \to" "R_t = 20\frac{20}{27}\varOmega"

a) we have "I_1 = 5A"       "U_1 = I_1*R_1 = 125V"

"U_{1,2,3,4} = U_1 = U_2=U_{3,4}=125V"  because of they are connected parallel

"I_{1,2,3,4} = \large\frac{U_{1,2,3,4}}{R_{1,2,3,4}}=\frac{125*11}{75}=\frac{55}{3}"

"I_{1,2,3,4} = I_5=I_{1,2,3,4,5}=\frac{55}{3}" because of they are connected in series

"U_5 = I_5*R_5=\frac{55}{3}*45=825V"

b) "U_{1,2,3,4,5} = U_{1,2,3,4}+U_{5}=825+125=1000V"

"U_t =U_{1,2,3,4,5} = U_6 = 1000V" because of they are connected parallel

c) "I_t = \large\frac{U_t}{T_t} = \frac{1000*27}{560} =" 48.2A